# On balanced Hermitian threefolds with parallel Bismut torsion

### v1

@@ Abstract.
We continue our study on Hermitian manifolds that are Bismut torsion parallel, or BTP for brevity, which means that the Bismut connection has parallel torsion tensor. For $n\geq 3$, BTP metrics can be balanced (and non-Kähler). In this paper, we give a classification of all compact, balanced BTP threefolds.
@@ 1. Introduction and statement of results
A Hermitian manifold $(M^{n},g)$ is said to be Bismut torsion parallel (or BTP for brevity), if its Bismut connection has parallel torsion tensor. Unless mentioned otherwise, we will always assume that $M^{n}$ is compact and $g$ is non-Kähler. Such manifolds form a relatively rich and highly interesting class of special Hermitian manifolds. It contains all Bismut Kähler-like (BKL) manifolds and all Vaisman manifolds as special examples. In complex dimension $2$, BTP $=$ BKL $=$ Vaisman by [34], and compact non-Kähler Vaisman surfaces are classified by Belgun in [8]. In complex dimensions $3$ or higher, BKL and Vaisman manifolds form proper, disjoint subsets of the set of non-balanced BTP manifolds. Also, starting in complex dimension $3$, there are examples of balanced (and non-Kähler) BTP manifolds, which tend to form a much smaller set compared with the non-balanced ones, as illustrated by complex nilmanifolds with BTP metrics.
@@
In [35], we studied the general properties of BTP manifolds and analyzed the structure of non-balanced BTP manifolds. In particular, the BTP condition, which means the parallelness of the Bismut torsion, is equivalent to some conditions involving the Bismut curvature tensor alone ([35, Theorem 1.1]). This is a rather distinctive property about the Bismut connection $\nabla^{b}$, and one certainly does not expect the same to hold for any other metric connection on $M^{n}$, as torsion and curvature are independent geometric invariants in general. In [35, Proposition 1.7], we also established the existence of a special kind of local unitary frames (called admissible frames) on any non-balanced BTP manifold, under which the Chern torsion components are particularly simple. This frame makes the analysis on non-balanced BTP manifolds more accessible.
@@
As mentioned before, in complex dimension $3$ or higher there are examples of compact balanced (but non-Kähler) BTP manifolds, although such manifolds tend to be fewer than non-balanced BTP ones. To justify this statement, note that Proposition 1.10 of [35] classified BTP metrics among all complex nilmanifolds with nilpotent $J$ in the sense of Cordero-Fernández-Gray-Ugarte [13]. It turns out that the majority of these BTP nilmanifolds are non-balanced, while only a small portion of them are balanced. For instance, when the complex dimension is $3$, only one of them (up to scaling of the metric) is balanced, while the non-balanced ones form a two-parameter family.
@@
For balanced BTP manifolds, we do not have the admissible frames to help us, which is why we are only able to deal with the $3$-dimensional situation in this paper. For $n=3$, Zhou and the second named author observed in [36] that any balanced Hermitian threefold $(M^{3},g)$ always admits a special type of local unitary frames (called special frames) where the Chern torsion take a relatively simple form. This technical tool enabled us to get the analysis started and to obtain characterization on all compact balanced BTP threefolds.
@@ 3. $(3c)$
In other words, a compact balanced BTP threefold is either Chern flat which is a compact quotient of the complex simple Lie group $SL(2,{\mathbb{C}})$ equipped with the standard metric, or the Wallach threefold, or of middle type. Up to a double cover if necessary, any middle type one is always primary and admits Vaisman companions, which are Vasiman threefolds sharing the same Bismut connection with the original primary threefold, thus having abelian Bismut holonomy group. See [20] for some very recent results on holonomy groups of Hermitian manifolds.
It follows directly from the proof of the main theorem that the first two cases above are actually BAS, which stands for Bismut Ambrose Singer. This notion was introduced in [21, 22] and it means that the Bismut connection $\nabla^{b}$ has parallel torsion and curvature, namely, $\nabla^{b}T^{b}=0$ and $\nabla^{b}R^{b}=0$. A classic theorem of Ambrose and Singer [6] (and its Hermitian version by Sekigawa [28]) says if a complete Riemannian (Hermitian) manifold admits a metric connection (Hermitian connection) which has parallel torsion and curvature, then the manifold is locally homogeneous, that is, its universal cover is a homogeneous Riemannian (Hermitian) manifold. In particular, BAS manifolds are always locally homogeneous. While, for the third case, namely, the middle type, it turns out that the entire Bismut (or Chern or Riemannian) curvature tensor is determined by two real-valued functions, $s$ and $\sigma_{2}$, which are the first and second elementary symmetric functions of the eigenvalues of the Bismut Ricci form. Of course $s$ is just the Bismut (Chern) scalar curvature. As observed in Proposition 1.7 of [24], this case will be BAS if and only if both $s$ and $\sigma_{2}$ are constants. See Lemma 5.6 for more details.
In Proposition 5.9, we classified all balanced BTP threefolds of middle type that are Lie-Hermitian manifolds, namely, compact quotients of unimodular Lie groups with left-invariant complex structure and compatible left-invariant metrics. They form two families $A_{s,t}$ and $B_{z,t}$, with $s,t\in{\mathbb{R}}$ and $z\in{\mathbb{C}}$, satisfying the followings
@@ 1. $(1)$
    $A_{0,0}=B_{0,0}=N^{3}$ is the only balanced BTP nilmanifold in dimension $3$.
@@ 5. $(5)$
Recall that a compact Hermitian manifold $(M^{n},g)$ is said to be CYT (which stands for Calabi-Yau with torsion), if the first Bismut Ricci curvature vanishes. It is said to be of Calabi-Yau type if its canonical line bundle is holomorphically trivial. Another result of this paper is the following
@@
Recall that Fino-Vezzoni Conjecture states that if a compact complex manifold admits a balanced metric and a pluriclosed metric, then it must admit a Kähler metric. Streets-Tian Conjecture states that any compact Hermitian-symplectic manifold must admit a Kähler metric. Both conjectures are known to be true for non-balanced BTP threefolds by [35, Corollary 1.19]. For balanced BTP threefolds, by Theorem 1.1 we just need to consider the middle type case, as the Fano case is already Kählerian while the Chern flat case is either known (for Streets-Tian conjecture in any dimension by [14]) or relatively easy to deduce (see the proof of Corollary 5.8).
@@
The Bismut connection [9] (also known as Strominger connection [30] in some literature) is the unique Hermitian connection with totally skew-symmetric torsion. So BTP manifolds are special examples of the much broader class of Riemannian manifolds admitting a metric connection with skew-symmetric and parallel torsion. There have been extensive studies on metric connections with parallel skew-symmetric torsion in the setting of Riemannian geometry or almost Hermitian geometry, by Friedrich, Agricola, Ferreira, Schoemann and others, see for example [1], [2], [3], [4], [5], [27] and the references therein. Here we carry out the investigation in the slightly more delicate Hermitian situation, where the emphasis is on complex structures, with the hope of understanding the Bismut geometry for complex manifolds.
The paper is organized as follows. In §2, we will discuss the types of the $B$ tensor of a non-Kähler balanced BTP threefold, prove Proposition 2.4, and also show that the $\mathrm{rank}\,B=3$ case of Theorem 1.1 leads to compact quotients of $SL(2,{\mathbb{C}})$. In §3, we will prove that the $\mathrm{rank}\,B=1$ case of Theorem 1.1 leads to the Wallach threefold. In §4, we will carry out the detailed computation which shows that the Wallach threefold is indeed balanced and BTP. In the Appendix, we will verify that the Wallach threefold has nonnegative bisectional curvature and positive holomorphic sectional curvature for its Chern connection, while its Riemannian connection has non-negative sectional curvature and constant Ricci curvature. In §5, we will prove the $\mbox{rank}\,B=2$ case of Theorem 1.1, prove Theorem 1.2, and classify Lie-Hermitian threefolds that are balanced BTP. Finally, in the last section, we will consider some generalization in higher dimensions, and prove Theorem 6.3.
@@ 2. Balanced BTP threefolds
We will follow the notations in [35]. Given a Hermitian manifold $(M^{n},g)$, denote by $\nabla^{c}$, $\nabla^{b}$ the Chern and Bismut connection, and by $T$, $T^{b}$ the Chern and Bismut torsion. Under a frame $e$ of type $(1,0)$ tangent vectors, let $T^{j}_{ik}$ be the components of the Chern torsion, namely,
@@
The BTP assumption on the metric $g$ means that $\nabla^{b}T^{b}=0$, which is equivalent to $\nabla^{b}T=0$.
Now let $(M^{3},g)$ be a compact non-Kähler balanced BTP manifold. We will start with a technical observation in [36, Proposition 2] which says that any balanced threefold always admits a particular type of unitary frame under which the Chern torsion takes a simple form. We include the proof here for readers’ convenience.
Lemma 2.1 ([36]).
Let $(M^{3},g)$ be a Hermitian threefold that is balanced. Then for any given $p\in M^{3}$, there exists a local unitary frame $e$ in a neighborhood of $p$ such that the only possibly non-zero components of the Chern torsion under $e$ are:
$T^{1}_{23}\geq T^{2}_{31}\geq T^{3}_{12}\geq 0.$
@@
Here we assume that $(ijk)$ and $(abc)$ are both cyclic permutations of $(123)$. Hence $\widetilde{A}$ is diagonal, which means that under the new unitary frame $\tilde{e}$ we have $\widetilde{T}^{r}_{rs}=0$ for any $1\leq r,s\leq 3$, and the only possibly non-zero torsion components are $\widetilde{T}^{i}_{jk}=-\widetilde{T}^{i}_{kj}$ where $(ijk)$ is any cyclic permutations of $(123)$. Next, let us write $a_{i}=\widetilde{T}^{i}_{jk}$. For each $1\leq i\leq 3$, define $\rho_{i}$ by
$\rho_{i}=\left\{\begin{array}[]{ll}\frac{a_{i}}{|a_{i}|},\ \ \ \mbox{if}\ \ a_{i}\neq 0,\\ \ \ 1,\ \ \ \ \mbox{if}\ \ a_{i}=0.\end{array}\right.$
Then for each $i$ we always have $|\rho_{i}|=1$ and $a_{i}=\rho_{i}|a_{i}|$. Consider the new unitary frame $e^{\prime}$ given by $e^{\prime}_{i}=(\overline{\rho}_{j}\overline{\rho}_{k})^{\frac{1}{2}}\,\tilde{e}_{i}$ where $(ijk)$ is a cyclic permutation of $(123)$. Then we still have $T^{\prime r}_{rs}=0$ for any $r$, $s$, while
@@
Interchange the orders of $e_{i}^{\prime}$ if necessary, we may assume that $a_{1}\geq a_{2}\geq a_{3}$, hence the frame $e^{\prime}$ satisfies the requirement of the lemma. ∎
@@
(2.1) $T^{1}_{1k}=T^{2}_{2k}=T^{3}_{3k}=0,\ \ \ \forall\ 1\leq k\leq 3\ \ \ \ \mbox{and}\ \ \ T^{1}_{23}\geq T^{2}_{31}\geq T^{3}_{12}\geq 0.$
Lemma 2.1 simply says that special frames exist around any given point on a balanced threefold. Now let $(M^{3},g)$ be a compact Hermitian threefold that is balanced and BTP. Let $e$ be a special frame and write
@@
Let $(M^{3},g)$ be a Hermitian threefold that is balanced and BTP. Suppose that $e$ is a special frame. Then the local non-negative functions $a_{1},a_{2},a_{3}$ given by (2.2) are global constants, and
[…]
@@
Let $B$ be the $2$-tensor on the manifold defined by $B_{i\overline{j}}=\sum_{r,s}T^{j}_{rs}\overline{T^{i}_{rs}}$ under any local unitary frame. It is globally-defined and Hermitian symmetric. Under a special frame $e$, we have
@@
By the BTP assumption, we have $\nabla^{b}B=0$, so the eigenvalues of $B$ are all global constants on $M$. Therefore each $a_{i}$ is a global constant.
@@
Denote by $\varphi$ the unitary coframe dual to the special frame $e$, and by $\theta$, $\tau$ the matrix of connection and column vector of torsion under $e$ for the Chern connection $\nabla^{c}$. Recall that the Chern torsion components $T^{j}_{ik}$ are defined by $T(e_{i},e_{k})=\sum_{j}T^{j}_{ik}e_{j}$, so we have $\tau_{j}=\frac{1}{2}\sum_{i,k}T^{j}_{ik}\,\varphi_{i}\wedge\varphi_{k}$.
Let $\gamma=\nabla^{b}-\nabla^{c}$ be the tensor, and for convenience we will also use the same letter to denote its matrix representation under $e$: $\gamma=\theta^{b}-\theta^{c}$. On any Hermitian manifold and under any unitary frame, it holds that $\gamma_{ij}=\sum_{k}(T^{j}_{ik}\varphi_{k}-\overline{T^{i}_{jk}}\overline{\varphi}_{k})$, Therefore in our case we have
@@
$0=d^{2}\varphi_{3}=a_{2}\,d(\varphi_{2}\overline{\varphi}_{1})-a_{1}\,d(\varphi_{1}\overline{\varphi}_{2})=2(a_{2}\overline{\psi}_{3}-a_{1}\psi_{3})\wedge(\varphi_{2}\,\overline{\varphi}_{2}-\varphi_{1}\,\overline{\varphi}_{1}).$
@@
$0=d^{2}\varphi_{3}=-a_{3}\,d(\varphi_{1}\varphi_{2})+a\,d(\varphi_{2}\overline{\varphi}_{1}-\varphi_{1}\overline{\varphi}_{2})=2a_{3}^{2}\varphi_{3}(\varphi_{1}\overline{\varphi}_{1}+\varphi_{2}\overline{\varphi}_{2}),$
@@
In the proof of Case 2 in Proposition 2.4, we already know that $(M^{3},g)$ is a compact Chern flat threefold. By [10], the universal cover of $M^{3}$ is a connected, simply-connected complex Lie group $G$, and the metric $\tilde{g}$ which is the lift of $g$ is left-invariant and is compatible with the complex structure of $G$. Let $\{\varepsilon_{1},\varepsilon_{2},\varepsilon_{3}\}$ be a left-invariant $\nabla^{c}$-parallel unitary frame of $G$. Denote by $T^{j}_{ik}$ the components of the Chern torsion under the frame $\varepsilon$, which are all constants. As $\tilde{g}$ is also non-Kähler balanced BTP, after a constant unitary transformation, we may assume that $\varepsilon$ is a special frame on $G$ and $a_{1}=a_{2}=a_{3}=a>0$ under this frame. Hence the Lie algebra of $G$ is ${\mathfrak{g}}={\mathbb{C}}\{\varepsilon_{1},\varepsilon_{2},\varepsilon_{3}\}$, satisfying
$[\varepsilon_{1},\varepsilon_{2}]=-a\varepsilon_{3},\ \ \ [\varepsilon_{2},\varepsilon_{3}]=-a\varepsilon_{1},\ \ \ [\varepsilon_{3},\varepsilon_{1}]=-a\varepsilon_{2}.$
@@
[…]

### v2

@@ Abstract.
We continue our study on Hermitian manifolds that are Bismut torsion parallel, or BTP for brevity, which means that the Bismut connection has parallel torsion tensor. For $n\geq 3$, BTP metrics can be balanced (and non-Kähler). In this paper, we give a detailed description to characterize all compact, balanced BTP threefolds.
@@ 1. Introduction and statement of results
A Hermitian manifold $(M^{n},g)$ is said to be Bismut torsion parallel (or BTP for brevity), if its Bismut connection has parallel torsion tensor. Unless mentioned otherwise, we will always assume that $M^{n}$ is compact and $g$ is non-Kähler. Such manifolds form a relatively rich and highly interesting class of special Hermitian manifolds. It contains all Bismut Kähler-like (or BKL) manifolds and all Vaisman manifolds as special examples. In complex dimension $2$, BTP == BKL == Vaisman by [ZZ23], and compact non-Kähler Vaisman surfaces are classified by Belgun in [Bel00]. In complex dimensions $3$ or higher, BKL and Vaisman manifolds form proper, disjoint subsets of the set of non-balanced BTP manifolds as shown in [YZZ23]. Also, starting in complex dimension $3$, there are examples of balanced (and non-Kähler) BTP manifolds, which tend to form a much smaller set compared with the non-balanced ones, as illustrated by complex nilmanifolds with BTP metrics.
@@
In [ZZ24+], we have studied the general properties of BTP manifolds and analyzed the structure of non-balanced BTP manifolds. In particular, the BTP condition, which means the parallelness of the Bismut torsion, is equivalent to some conditions involving the Bismut curvature tensor alone, by [ZZ24+, Theorem 1.1]. This is a rather distinctive property about the Bismut connection $\nabla^{b}$, and one certainly does not expect the same to hold for any other metric connection on $M^{n}$, as torsion and curvature are independent geometric invariants in general. In [ZZ24+, Proposition 1.7], we also established the existence of a special kind of local unitary frames (called admissible frames) on any non-balanced BTP manifold, under which the Chern torsion components are particularly simple. This frame makes the analysis on non-balanced BTP manifolds more accessible, so we get a structure theorem of such manifolds of complex dimension $3$ in [ZZ24+, Theorem 1.16].
@@
As mentioned before, in complex dimension $3$ or higher there are examples of compact balanced (but non-Kähler) BTP manifolds, although such manifolds tend to be fewer than non-balanced BTP ones. To justify this statement, note that Proposition 1.10 of [ZZ24+] classified BTP metrics among all complex nilmanifolds with nilpotent $J$ in the sense of Cordero-Fernández-Gray-Ugarte [CFGU00]. It turns out that the majority of these BTP nilmanifolds are non-balanced, while only a small portion of them are balanced. For instance, when the complex dimension is $3$, only one of them (up to scaling of the metric) is balanced, while the non-balanced ones form a two-parameter family.
@@
For balanced BTP manifolds, we do not have the admissible frames to help us, which is why we are only able to deal with the $3$-dimensional situation in this paper. For $n=3$, Zhou and the second named author observed in [ZhZ22] that any balanced Hermitian threefold $(M^{3},g)$ always admits a special type of local unitary frames (called special frames) where the Chern torsion take a relatively simple form. This technical tool enabled us to get the analysis started and to obtain characterization on all compact balanced BTP threefolds.
@@ 3. $(3c)$
In other words, a compact balanced BTP threefold is either Chern flat which is a compact quotient of the complex simple Lie group $SL(2,{\mathbb{C}})$ equipped with the standard metric, or the Wallach threefold, or of middle type. Up to a double cover if necessary, any middle type one is always primary and admits Vaisman companions, which are Vasiman threefolds sharing the same Bismut connection with the original primary threefold, thus having abelian Bismut holonomy group. See [Ni25+] for some very recent results on holonomy groups of Hermitian manifolds.
It follows directly from the proof of the main theorem that the first two cases above are actually BAS, which stands for Bismut Ambrose Singer. This notion was introduced in [NZ23, NZ23+] and it means that the Bismut connection $\nabla^{b}$ has parallel torsion and curvature, namely, $\nabla^{b}T^{b}=0$ and $\nabla^{b}R^{b}=0$. A classic theorem of Ambrose and Singer [AS58] (and its Hermitian version by Sekigawa [Sek78]) says if a complete Riemannian (Hermitian) manifold admits a metric connection (Hermitian connection) which has parallel torsion and curvature, then the manifold is locally homogeneous, that is, its universal cover is a homogeneous Riemannian (Hermitian) manifold. In particular, BAS manifolds are always locally homogeneous. While, for the third case, namely, the middle type, it turns out that the entire Bismut (or Chern or Riemannian) curvature tensor is determined by two real-valued functions, $s$ and $\sigma_{2}$, which are the first and second elementary symmetric functions of the eigenvalues of the Bismut Ricci form. Of course $s$ is just the Bismut (Chern) scalar curvature. As observed in Proposition 1.7 of [PZ25], this case will be BAS if and only if both $s$ and $\sigma_{2}$ are constants. See Lemma 5.6 for more details.
In Proposition 5.9 and the content after it, we classified $6$-dimensional unimodular Lie algebras which admits a Hermitian structure that is balanced BTP of the middle type. We discussed the isomorphic classes of these Lie algebras and the existence of uniform lattices in the corresponding simply-connected Lie groups. They form two families $A_{s,t}$ and $B_{z,t}$ for $s,t\in{\mathbb{R}}$ and $z\in{\mathbb{C}}$, satisfying the followings
@@ 1. $(1)$
    $A_{0,0}=B_{0,0}=N^{3}$ determines the only balanced BTP nilmanifold in dimension $3$.
@@ 5. $(5)$
Recall that a compact Hermitian manifold $(M^{n},g)$ is said to be CYT, which stands for Calabi-Yau with torsion, if the first Bismut Ricci curvature vanishes. It is said to be of Calabi-Yau type if its canonical line bundle is holomorphically trivial. Another result of this paper is the following
@@
Recall that Fino-Vezzoni Conjecture states that if a compact complex manifold admits a balanced metric and a pluriclosed metric, then it must admit a Kähler metric. Streets-Tian Conjecture states that any compact Hermitian-symplectic manifold must admit a Kähler metric. Both conjectures are known to be true for non-balanced BTP threefolds by [ZZ24+, Corollary 1.19]. For balanced BTP threefolds, by Theorem 1.1 we just need to consider the middle type case, as the Fano case is already Kählerian while the Chern flat case is either known (for Streets-Tian conjecture in any dimension by [1]) or relatively easy to deduce as in the proof of Corollary 5.8.
@@
The Bismut connection [Bis89], also known as Strominger connection [Str86] in the literature, is the unique Hermitian connection with totally skew-symmetric torsion. So BTP manifolds are special examples of the much broader class of Riemannian manifolds admitting a metric connection with skew-symmetric and parallel torsion. There have been extensive studies on the latter in the setting of Riemannian geometry or almost Hermitian geometry with an incomplete list [AF04], [AF14], [AFF15], [AFS05], [AFK08], [CMS21] [Sch07] and the references therein. Here we carry out the investigation in the slightly more delicate Hermitian situation, where the emphasis is on complex structures, with the hope of understanding the Bismut geometry for complex manifolds.
The paper is organized as follows. In §2, we will discuss the types of the $B$ tensor of a non-Kähler balanced BTP threefold, prove Proposition 2.4, and also show that the $\mathrm{rank}\,B=3$ case of Theorem 1.1 leads to compact quotients of $SL(2,{\mathbb{C}})$. In §3, we will prove that the $\mathrm{rank}\,B=1$ case of Theorem 1.1 leads to the Wallach threefold. In §4, we will carry out the detailed computation which shows that the Wallach threefold is indeed balanced and BTP. In the Appendix, we will verify that the Wallach threefold has nonnegative bisectional curvature and positive holomorphic sectional curvature for its Chern connection, while its Riemannian connection has non-negative sectional curvature and constant Ricci curvature. In §5, we will prove the $\mbox{rank}\,B=2$ case of Theorem 1.1, prove Theorem 1.2, and classify Hermitian unimodular Lie algebras of complex dimension $3$ which are balanced BTP of the middle type. Finally, in the last section, we will consider some generalization in higher dimensions, and prove Theorem 6.3.
@@ 2. Balanced BTP threefolds
We will follow the notations in [ZZ24+] throughout this paper. Given a Hermitian manifold $(M^{n},g)$, denote by $\nabla^{c}$, $\nabla^{b}$ the Chern and Bismut connection, and by $T$, $T^{b}$ the Chern and Bismut torsion. Under a frame $e$ of type $(1,0)$ tangent vectors, let $T^{j}_{ik}$ be the components of the Chern torsion, namely,
@@
The BTP assumption on the metric $g$ means that $\nabla^{b}T^{b}=0$, which is equivalent to $\nabla^{b}T=0$. This has been shown in [ZZ23, the proof of Proposition 1]
Now let $(M^{3},g)$ be a compact non-Kähler balanced BTP manifold. We will start with a technical observation in [ZhZ22, Proposition 2] which says that any balanced threefold always admits a particular type of unitary frame under which the Chern torsion takes a simple form. We include the proof here for readers’ convenience. Let us denote by $B$ the $2$-tensor on the manifold defined by $B_{i\overline{j}}=\sum_{r,s}T^{j}_{rs}\overline{T^{i}_{rs}}$ under any local unitary frame. It is globally-defined and Hermitian symmetric.
Lemma 2.1 ([ZhZ22]).
Let $(M^{3},g)$ be a Hermitian threefold that is balanced. Then for any given $p\in M^{3}$, there exists a local unitary frame $e$ in a neighborhood of $p$ such that the $T^{i}_{ij}=0$ for $1\leq i,j\leq 3$ under $e$. Furthermore, if the rank of the $B$ tensor is constant in the neighborhood, then we can choose a local unitary frame $e$ so that $T^{i}_{ij}=0$ for $1\leq i,j\leq 3$ and
$T^{1}_{23}\geq 0,\ T^{2}_{31}\geq 0,\ T^{3}_{12}\geq 0.$
@@
Here we assumed that $(ijk)$ and $(abc)$ are both cyclic permutations of $(123)$. Hence $\widetilde{A}$ is diagonal, which means that under the new unitary frame $\tilde{e}$ we have $\widetilde{T}^{s}_{st}=0$ for any $1\leq s,t\leq 3$, and the only possibly non-zero torsion components are $\widetilde{T}^{i}_{jk}=-\widetilde{T}^{i}_{kj}$ where $(ijk)$ is any cyclic permutations of $(123)$. Let us write $a_{i}=\widetilde{T}^{i}_{jk}$. The $B$ tensor under $\tilde{e}$ now takes the form
$B=2\begin{bmatrix}|a_{1}|^{2}&0&0\\ 0&|a_{2}|^{2}&0\\ 0&0&|a_{3}|^{2}\end{bmatrix}\!\!$.
Denote by $r$ the constant rank of $B$ in the neighborhood. At $p$, exactly $r$ of those $a_{i}$ are non-zero. Without loss of generality, we may assume that $a_{1}\cdots a_{r}|_{p}\neq 0$. Let $V$ be a possibly smaller neighborhood of $p$ in which $a_{1}\cdots a_{r}\neq 0$. Since the rank of $B$ is constantly $r$, within $V$ we must have $a_{r+1}=\cdots=a_{3}=0$. Define smooth functions in $V$
$\rho_{i}=\frac{a_{i}}{|a_{i}|},\quad\forall\ 1\leq i\leq r\quad\mbox{and}\quad\rho_{i}=1,\quad\forall\ r\!+\!1\leq i\leq 3.$
Consider a new unitary frame $e^{\prime}$ in $V$ given by $e^{\prime}_{i}=(\overline{\rho}_{j}\overline{\rho}_{k})^{\frac{1}{2}}\,\tilde{e}_{i}$ where $(ijk)$ is a cyclic permutation of $(123)$. Then we still have $T^{\prime s}_{st}=0$ for any $s$, $t$, while
@@
Hence the frame $e^{\prime}$ satisfies the requirement of the lemma. ∎
@@
(2.1) $T^{1}_{1k}=T^{2}_{2k}=T^{3}_{3k}=0,\quad\forall\ 1\leq k\leq 3\quad\mbox{and}\quad T^{1}_{23}\geq 0,\ T^{2}_{31}\geq 0,\ T^{3}_{12}\geq 0.$
Lemma 2.1 simply says that special frames exist in a neighborhood of any given point in a balanced threefold wherever the $B$ tensor has constant rank. Now let $(M^{3},g)$ be a compact Hermitian threefold that is balanced and BTP. Then the condition $\nabla^{b}B=0$ guarantees that $B$ has constant rank over the entire $M$ as the eigenfunctions of $B$ are constants, so there always exist special frames everywhere. Let $e$ be a special frame and write
@@
Let $(M^{3},g)$ be a Hermitian threefold that is balanced and BTP. Suppose that $e$ is a special frame. Then the local non-negative functions $a_{1},a_{2},a_{3}$ given by (2.2) are global constants and we may assume $a_{1}\geq a_{2}\geq a_{3}$ after a signed permutation of $e$. Furthermore, it holds that
(2.3) $\left\{\begin{split}a_{1}(\theta^{b}_{22}+\theta^{b}_{33}-\theta^{b}_{11})\ =\ 0,\hskip 133.72786pt\\ a_{2}(\theta^{b}_{11}+\theta^{b}_{33}-\theta^{b}_{22})\ =\ 0,\hskip 133.72786pt\\ a_{3}(\theta^{b}_{11}+\theta^{b}_{22}-\theta^{b}_{33})\ =\ 0,\hskip 133.72786pt\\ a_{1}\theta^{b}_{12}+a_{2}\theta^{b}_{21}\ =\ a_{1}\theta^{b}_{13}+a_{3}\theta^{b}_{31}\ =\ a_{2}\theta^{b}_{23}+a_{3}\theta^{b}_{32}\ =\ 0,\end{split}\right$.
@@
The $B$ tensor under a special frame $e$ takes the form
@@
By the BTP assumption, we have $\nabla^{b}B=0$, so the eigenfunctions of $B$ are all global constants on $M$, which turn out to be $2a_{1}^{2}$, $2a_{2}^{2}$ and $2a_{3}^{2}$. Therefore each $a_{i}$ is a global constant. If these non-negative constants $a_{1},a_{2},a_{3}$ are not in non-increasing order, say for instance $a_{3}\geq a_{1}\geq a_{2}$, then by using the special frame $(e_{3},e_{1},e_{2})$ instead, we get them into the non-increasing order. For another example, if $a_{3}\geq a_{2}\geq a_{1}$, then we may use the special frame $(-e_{3},e_{2},e_{1})$ to achieve the requirement.
@@
Denote by $\varphi$ the unitary coframe dual to the special frame $e$, and by $\theta$, $\tau$ the matrix of connection and column vector of torsion under $e$ for the Chern connection $\nabla^{c}$. As the Chern torsion components $T^{j}_{ik}$ are defined by $T(e_{i},e_{k})=\sum_{j}T^{j}_{ik}e_{j}$, so we have $\tau_{j}=\frac{1}{2}\sum_{i,k}T^{j}_{ik}\,\varphi_{i}\wedge\varphi_{k}$ by the structure equations and Bianchi identities set up in [ZZ24+, Section 2].
Let $\gamma=\nabla^{b}-\nabla^{c}$ be the $(2,1)$-tensor introduced in [YZ18], and for convenience we will also use the same letter to denote its matrix representation under $e$, that is, $\gamma=\theta^{b}-\theta^{c}$. On the Hermitian manifold $(M^{n},g)$ and under any unitary frame, it holds that $\gamma_{ij}=\sum_{k}(T^{j}_{ik}\varphi_{k}-\overline{T^{i}_{jk}}\overline{\varphi}_{k})$ as shown in [YZ18]. These have also been given in [ZZ24+, Section 2]. Therefore in our case we have
@@
$0=d^{2}\varphi_{3}=a_{2}\,d(\varphi_{2}\overline{\varphi}_{1})-a_{1}\,d(\varphi_{1}\overline{\varphi}_{2})=(a_{2}\overline{\psi}_{3}-a_{1}\psi_{3})\wedge(\varphi_{2}\,\overline{\varphi}_{2}-\varphi_{1}\,\overline{\varphi}_{1}).$
@@
$0=d^{2}\varphi_{3}=-a_{3}\,d(\varphi_{1}\varphi_{2})+a\,d(\varphi_{2}\overline{\varphi}_{1}-\varphi_{1}\overline{\varphi}_{2})=a_{3}^{2}\varphi_{3}(\varphi_{1}\overline{\varphi}_{1}+\varphi_{2}\overline{\varphi}_{2}),$
@@
In the proof of Case 2 in Proposition 2.4, we already know that $(M^{3},g)$ is a compact Chern flat threefold. By [Boo58], the universal cover of $M^{3}$ is a connected, simply-connected complex Lie group $G$, and the metric $\tilde{g}$ which is the lift of $g$ is left-invariant and is compatible with the complex structure of $G$. Let $\{\varepsilon_{1},\varepsilon_{2},\varepsilon_{3}\}$ be a left-invariant $\nabla^{c}$-parallel unitary frame of $G$. Denote by $T^{j}_{ik}$ the components of the Chern torsion under the frame $\varepsilon$, which are all constants. As $\tilde{g}$ is also non-Kähler balanced BTP, after a constant unitary transformation, we may assume that $\varepsilon$ is a special frame on $G$ and $a_{1}=a_{2}=a_{3}=a>0$ under this frame. Hence the Lie algebra of $G$ is ${\mathfrak{g}}={\mathbb{C}}\{\varepsilon_{1},\varepsilon_{2},\varepsilon_{3}\}$, satisfying
$\varepsilon_{1},\varepsilon_{2}]=-a\varepsilon_{3},\ \ \ [\varepsilon_{2},\varepsilon_{3}]=-a\varepsilon_{1},\ \ \ [\varepsilon_{3},\varepsilon_{1}]=-a\varepsilon_{2}.$
@@
$\sum_{\alpha}\overline{P_{i\alpha}}S_{j\alpha}=\delta_{ij},\ \ \ \forall\ 1\leq i,j\leq 3.$